Ticket Number: HEA-NOC/20011002-01 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> CMOD
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20011003 08.00 UTC+1
Ticket Closed: 20011019 15:38:07 UTC Problem End: 20011019 16:35 UTC+1
Eirc:om are to upgrade the CMOD line (DS 0838705), from 512K to 2M.
There will be a brief outage during this upgrade (the length of which
and the time it will start has yet to be confirmed) which will affect
all traffic to and from CMOD.
Am awaiting confirmation from Eircom detailing when this outage will
occur, and the expected length of the outage.
OMG 15.22 - CMOD have confirmed that this outage does not suit them,
and they are arranging a rescheduling with Eircom for this upgrade. This ticket will be updated once we have details of the new date for
20011009 14.43 UTC+1 OMG - This upgrade has been provisionally re-
scheduled for Wednesday 20011017 at 09.00 UTC+1. The upgrade is on the
existing line, and there shouldn't be an outage of more than 1 hour.
Another ticket will be posted when we have confirmation from Eircom.
20011017 SY 12:24 UTC+1 CMOD down on our monitoring tools form 10:43, Eircom have upgraded the line without prior notification Eircom say the line is clear at both ends we tried resetting the link simultaneously at both ends without success. Eddie is going to contact Lancomms for advice
20011017 SY 14:31 UTC+1 CMOD up Eircom put local loops on the circuit and reversed the timing on the interface on the Martis node in Brooklawn House, packets are flowing over the circuit without errors will continue to monitor.
20011019 SY 16:30 UTC+1 CMOD cicuit has no errors since 14:31 on 17th Oct, left a message with Eddie McGinn on his answerphone if he wants to get back to [log in to unmask] with an other issues regarding the upgrade, ticket closed.
Time to Fix: